# General theory for extended-range percolation on simple and multiplex networks

### v1

@@ I Introduction
Recently there is an increasing interest in percolation problem that do not require the traditional notion of connectivity. In addition to ERP, several other percolation models have been recently proposed in which percolation is defined through properties of the shortest paths of the network or through generalized notions of connected components. These include concurrence and $\alpha$-percolation [23, 4], motivated by quantum communication, shortest path percolation, [24] motivated by transportation networks, and no-exclaves percolation [25] motivated by the need to enhance network robustness. Moreover models assuming a nonlocal definition of connectivity are raising interest in the last few years, also beyond the theory of percolation with many applications including notably epidemic spreading [26, 27]. Thus a general theory of ERP on simple and multilayer networks might inspire further research in this direction.
@@
In ERP nodes can be trusted with probability $p$ and untrusted with probability $1-p$. Untrusted nodes can be involved in the communication between trusted nodes if they lie on a path between trusted nodes of length at most $R$. The order parameter $P^{\infty}$ of ERP is given by the size of the extended-range giant component (ERGC) formed by trusted nodes connected by paths of trusted and facilitator nodes. A facilitator is an untrusted node lying on the shortest path of length at most $R$ between two trusted nodes. Recently the exact solution of extended-range percolation on uncorrelated random graphs with given degree distribution was found for $R$ up to $6$ [15]. Here we provide a general theory of ERP for any arbitrary $R$. This theory is based on a message passing (MP) approach which leads to MP equations that can be used to study the ERP on any arbitrary locally tree-like network. The message passing approach  [28, 14, 29] defines a fundamental distributed computation that applies to a large variety of critical phenomena and dynamical process on networks including percolation, network control, and optimization problems [30, 31, 32, 33, 34, 35]. For standard percolation the MP predicts the size of the giant component on arbitrary network topology as long as it is locally tree-like; extensions beyond this approximation are a topic of active research activity [36, 37]. Here we fully develop a message passing theory of ERP that allows to predict the size of the ERGC for any value of $R$ and for any arbitrary network that is locally tree-like. In this work we also use this theory to investigate ERP in random graphs with given degree distribution. We stress that, although the formalism and approach developed here is markedly distinct from the one formulated in Ref. [15], the equations that we obtain for arbitrary $R$ are equivalent to the ones obtained by the previous work for $R$ up to $6$.
@@
We provide an exact solution of both versions of MERP on uncorrelated random multiplex networks with arbitrary degree distributions. Our analysis of the critical phenomena of MERP is based upon analyticalpredictions of the phase diagram, supported by extensive Monte Carlo simulations. We demonstrate an important interplay between the effect of interdependecies and of the extended-range mechanism. In particular, the study of version $V=2$ of MERP shows a highly nontrivial phenomenology with a reentrant discontinuous hybrid phase transition, observed in multiplex networks where the MCGC does not exist.
@@ II Extended-range percolation and message passing approach
Consider a network $\mathcal{G}=(V,E)$ of $N=|V|$ nodes where nodes are either untrusted or trusted. Communication between two trusted nodes takes place only if there is at least one walk connecting them including trusted and untrusted nodes and having at most $R-1$ consecutive untrusted nodes. Note that we require the existence of a walk and not a path. Thus we allow self-overlaps. For instance, nodes $3$ and $8$ in Fig. 1(a) can communicate for $R=3$ even if the (shortest) path $\{3\to 4\to 6\to 7\to 8\}$ connecting them contains three consecutive untrusted nodes, thanks to the existence of the walk $\{3,4,6,5,6,7,8\}$ having at most two consecutive untrusted nodes thanks to the presence of the trusted node $5$.
Given this way to communicate, we define the extended-range giant component (ERGC) of a simple network in terms of the trusted and the untrusted facilitator nodes. Untrusted facilitator nodes or simply facilitators nodes are untrusted nodes that can still ensure communications between trusted nodes if they lie on a path of length at most $R$ between two trusted nodes. For instance, nodes $4,6,7$ in Fig. 1 are facilitators for $R=3$. The extended-range giant component (ERGC) is formed by trusted nodes connected by paths formed exclusively by trusted or untrusted facilitator nodes. This is equivalent to say that any pair of trusted nodes in the ERGC is connected by at least a walk formed by trusted and untrusted nodes including at most $R-1$ consecutive untrusted nodes (facilitators). Additionally the ERGC includes all the untrusted nodes at distance less than $R$ from trusted nodes in the ERGC. For a schematic representation of the ERGCs see Fig. 1.
@@
Setting $R=1$, we recover the MP equations for standard percolation, [14].
@@ IV.1 Theoretical framework
${S}_{1}^{[\alpha]}$ $=$ $p\left[1-G_{1}^{[\alpha]}\left(1-\sum_{1\leq r\leq R}S_{r}^{[\alpha]}\right)\right]\left\{\prod_{\beta\neq\alpha}\left[1-G_{0}^{[\beta]}\left(1-\sum_{1\leq r\leq R}S_{r}^{[\beta]}\right)\right]\right\},$
${W}_{1}^{[\alpha]}$ $=$ $p\left[G_{1}^{[\alpha]}\left(1-\sum_{1\leq r\leq R}S_{r}^{[\alpha]}\right)\right]\left\{\prod_{\beta\neq\alpha}\left[1-G_{0}^{[\beta]}\left(1-\sum_{1\leq r\leq R}S_{r}^{[\beta]}\right)\right]\right\}.$ (15)
@@ IV.2 Phenomenology of the MERP for $R\in\{2,3,4\}$ on Poisson multiplex networks
Let us indicate with $\bar{c}$ the minimum average degree required to observe a MCGC (or equivalently a MERGC with $R=1$). Its value for $M=2$ is $\bar{c}=2.45541\ldots$. For version $V=1$ of MERP $\bar{c}$ also indicates the minimum average degree required to observe a MERGC independently of the value of $R\geq 1$. In version $V=1$ we observe a single hybrid discontinuous phase transition at the emergence of the MERGC.

### v2

@@ I Introduction
There is an increasing interest in percolation problems that do not require the traditional notion of connectivity. In addition to ERP, several other percolation models have been recently proposed in which percolation is defined through properties of the shortest paths of the network or through generalized notions of connected components. These include concurrence and $\alpha$-percolation [23, 4], motivated by quantum communication, shortest path percolation [24], motivated by transportation networks, no-exclaves percolation [25], motivated by the need to enhance network robustness , and color-avoiding percolation [26, 27], to reproduce nonuniform vulnerability of nodes to attack or failure. Moreover models assuming a nonlocal definition of connectivity are raising interest in the last few years, also beyond the theory of percolation with many applications including notably epidemic spreading [28, 29]. Thus a general theory of ERP on simple and multilayer networks might inspire further research in this direction.
In ERP nodes can be trusted with probability $p$ and untrusted with probability $1-p$. Untrusted nodes can be involved in the communication between trusted nodes if they lie on a path between trusted nodes of length at most $R$. We call these nodes facilitator nodes. The extended-range giant component (ERGC) is formed by all the trusted nodes connected by paths of trusted and facilitator nodes and a by all the untrusted nodes at distance less than $R$ from these trusted nodes. Recently the exact solution of extended-range percolation on uncorrelated random graphs with given degree distribution was found for $R$ up to $6$ [15]. However this solution relies on involved combinatorial definitions which impede a straightforward extension of the approach to treat larger values of $R$ as well as to address generalized percolation problems on multilayer or higher-order networks.
@@
Here we provide an general theory of ERP for any arbitrary $R$ that is based on a message passing (MP) approach where the messages have a simple and transparent combinatorial interpretation allowing a deeper theoretical understanding of the model as well as an easier generalization to multilayer networks. The message passing approach  [30, 14, 31] defines a fundamental distributed computation that applies to a large variety of critical phenomena and dynamical processes on networks including percolation, network control, and optimization problems [32, 33, 34, 35, 36, 37]. For standard percolation the MP predicts the size of the giant component on arbitrary network topology as long as it is locally tree-like; extensions beyond this approximation are a topic of active research [38, 39]. Here we fully develop a message passing theory of ERP that allows us to predict the size of the ERGC for any value of $R$ and for any arbitrary network that is locally tree-like. In this work we also use this theory to investigate ERP in random graphs with given degree distribution. We stress that, although the formalism and approach developed here is markedly distinct from the one formulated in Ref. [15], the equations that we obtain for arbitrary $R$ are equivalent to the ones obtained by the previous work for $R$ up to $6$.
@@
We provide an exact solution of both versions of MERP on uncorrelated random multiplex networks with arbitrary degree distributions. Our analysis of the critical phenomena of MERP is based upon analytical predictions of the phase diagram, supported by extensive Monte Carlo simulations. We reveal the important interplay between the effect of interdependecies and of the extended-range mechanism. In particular, the study of version $V=2$ of MERP shows a highly nontrivial phenomenology with a reentrant discontinuous hybrid phase transition, observed in multiplex networks where the MCGC does not exist.
@@ II Extended-range percolation and message passing approach
Consider a network $\mathcal{G}=(V,E)$ of $N=|V|$ nodes where nodes are either untrusted or trusted. Communication between two trusted nodes takes place only if there is at least one walk connecting them including trusted and untrusted nodes and having at most $R-1$ consecutive untrusted nodes. Note that we require the existence of a walk and not a path. For instance, nodes $3$ and $8$ in Fig. 1(a) can communicate for $R=3$ even if the (shortest) path $\{3\to 4\to 6\to 7\to 8\}$ connecting them contains three consecutive untrusted nodes, thanks to the existence of the walk $\{3,4,6,5,6,7,8\}$ having at most two consecutive untrusted nodes because of the presence of the trusted node $5$.
Strictly speaking, communication is among trusted nodes, but it is clear that untrusted nodes play a crucial role in this process. In particular, untrusted nodes lying on paths between two trusted nodes of length at most $R$. We call such untrusted nodes facilitators, because they are allowing for the communication of distant trusted nodes. For instance, node $2$ is a facilitator for $R=2$ and $R=3$ in Fig. 1, nodes $4,6,7$ are facilitators for $R=3$ in Fig. 1(a) but not in (b). Given this way to communicate, we define the extended-range giant connected component (ERGC) as the giant component formed by trusted nodes connected by paths including exclusively trusted or untrusted facilitator nodes and by all the untrusted nodes at distance less than $R$ from these trusted nodes. This is equivalent to say that any pair of trusted nodes in the ERGC is connected by at least a walk formed by trusted and untrusted nodes including at most $R-1$ consecutive untrusted nodes (facilitators). Moreover the untrusted nodes in the ERGC include all the facilitator nodes and all the untrusted nodes that are not facilitators but are at distance less than $R$ from the trusted nodes in the ERGC. For a schematic representation of the ERGCs see Fig. 1.
The algorithm to evaluate the ERGC is a generalization of the Depth-First Search algorithm. We assign to each node $i$ of the network a pair of variables $(q_{i},r_{i})$. Initially no node is assigned to a cluster, i.e. every node has $q_{i}=0$. Moreover all trusted nodes have $r_{i}=0$ and all untrusted nodes have $r_{i}=R$. At the end of the algorithm $q_{i}=q\in\{1,2,\ldots,Q\}$ indicates to which of the $Q$ clusters node $i$ belongs and $r_{i}$ indicates the distance (if smaller than $R$) from the closer trusted node. The algorithm for detecting the ERGC is the recursive Extended-Range Depth First Search ERDFS algorithm. Starting from a an arbitrary trusted (unvisited) seed node for cluster $q$ (initially taken to be $q=1$), we set $q_{i}=q$ and $r_{i}=0$ and we call the ERDFS algorithm.
The recursive ERDFS algorithm is defined as follows. Starting from a node $n$ with $q_{n}=q$ and $r_{n}=r$, ERDFS implements (1) and (2) defined as:
  • (1)
    If $r<R$ the algorithm visits all the neighbour trusted nodes $m$ unvisited previously. For each of these nodes the algorithm sets $r_{m}=0$ and $q_{m}=q$, and the ERDFS is recursively iterated.
  • (2)
    If $r<R-1$ the algorithm visits all the neighbour untrusted nodes $m$ with $r_{m}>r+1$. For each of these nodes the algorithm sets $r_{m}=r+1$ and $q_{m}=q$, and the ERDFS is recursively iterated.
When there are no more nodes to visit, the ERDFS stops and all the trusted and untrusted nodes with $q_{i}=q$ form the ERP cluster $q$. If there are unvisited trusted nodes we choose a new unvisited trusted seed node for cluster $q+1$ and iterate the procedure until and all the trusted nodes are assigned to a cluster and we have $Q$ clusters (where $Q$ is determined by the algorithm). The ERGC is the giant cluster $q=q^{\star}$ formed by the trusted nodes with $r_{i}=0$ and the untrusted nodes with $r_{i}<R$ all having $q_{i}=q^{\star}$.
@@
The overall fraction of nodes in the ERGC is the sum between $P^{\infty}$ and $U^{\infty}$. Setting $R=1$, we recover the MP equations for standard percolation, [14].
@@ IV.1 Theoretical framework
${S}_{1}^{[\alpha]}$ $=$ $p\left[1-G_{1}^{[\alpha]}\left(1-\sum_{1\leq r\leq R}S_{r}^{[\alpha]}\right)\right]$
$\left\{\prod_{\beta\neq\alpha}\left[1-G_{0}^{[\beta]}\left(1-\sum_{1\leq r\leq R}S_{r}^{[\beta]}\right)\right]\right\},$
${W}_{1}^{[\alpha]}$ $=$ $p\left[G_{1}^{[\alpha]}\left(1-\sum_{1\leq r\leq R}S_{r}^{[\alpha]}\right)\right]$ (15)
$\left\{\prod_{\beta\neq\alpha}\left[1-G_{0}^{[\beta]}\left(1-\sum_{1\leq r\leq R}S_{r}^{[\beta]}\right)\right]\right\}.$
@@ IV.2 Phenomenology of the MERP for $R\in\{2,3,4\}$ on Poisson multiplex networks
Let us indicate with $\bar{c}$ the minimum average degree required to observe a MCGC (or equivalently a MERGC with $R=1$). Its value for $M=2$ is $\bar{c}=2.45541\ldots$. For version $V=1$ of MERP $\bar{c}$ also indicates the minimum average degree required to observe a MERGC independently of the value of $R\geq 1$. In version $V=1$ and as long as $c>\bar{c}$ also in version $V=2$ we observe a single hybrid discontinuous phase transition at the emergence of the MERGC (see Fig. 5).